directions; use inductive reasoning to describe each pattern. then find the next two numbers in each pattern.
problem;
4, 6, 9, 13.5....
what are the next two numbers and how did you get it?

Answer by LtAurora(115)

The pattern is the number + 1/2 the number = the next number.
So:
4+2=6
6+3=9
9+4.5=13.5
This could also be written as: x+.5*x=Next number
So, the next two numbers are:
13.5+6.75=20.25
20.25+10.125=30.375